Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, bersanetti, ruggi - 23:24 Wednesday 05 May 2021 (51669) Print this report
ISC shift: CARM offset reduction

The goal of this shift was to proceed on the reduction of the CARM offset.
At the beginning of the shift we noticed glitches on B7 and B8 DC. As suggested by Chiummo, we tilted the SR mirror, starting from its parking position, by +30 urad on TX. This solved the issue (Fig.1).

We moved on by trying to lock the green on the arms and proceed to the CARM/DARM basis, but the North arm was too unstable to accomplish this, as it unlocked and re-locked repeatedly. This might have been caused by intense sea activity (which remained an issue for the whole shift) and by an oscillation of the SNEB bench, partially solved by an intervention by Ruggi and Vardaro.

We decided to focus our attention on the CITF standalone, with the idea of integrating the hand-off to the 3f signals for the alignment in the same Metatron node that manages this step for the longitudinal DoFs. We took the occasion to double check the condition of the CITF after the work on the IMC of last week. In particular:

  • we checked the demod phase for MICH and SRCL for the 1f signals, which was not completed yesterday, using the same methodology tried to reduce the MICH/PRCL coherence.
  • we recovered also the 3f error signals by tuning the demodulation phases (using the UGF lines as noise injections), and by re-measuring the ratios between 1f/3f signals.
  • we managed to perform the hand-off smoothly for all of the three loops.

We repeated the hand-off to the 3f signals for the PR alignment loop, and worked reliably. We tried to engage the (1f) drift control for the BS, after tuning again the demod phase for its error signal, but this loop proved once again to be unreliable and was ultimately disengaged for the rest of the shift.

After these tunings, the lock of the CITF and the hand-off to the 3f signals is reliable and robust.

We integrated all the mentioned changes in the automation, namely:

  • updated the demod phase for MICH and SRCL 1f
  • updated the ratios for the 1f/3f hand-off
  • updated MICH and PRCL gains
  • increased by 50% the PR_TX loop gain
  • changed the phase for B1p_QD2_6MHz for the BS, but the drift control still doesn't work well
  • we un-commented the gains for the BS drift control loop, since this is hardly the best way to leave it disengaged. Instead, we commented two couples of flags inside the python code (lines 527-528 and 541-542)
  • we added the parameters for the PR alignment loop with the 3f in the 3f section, and the hand-off is integrated in the automation in the analogous step for the longitudinal DoFs.
  • in the DOWN state for the DRMI, we now removed all the possible sensors for the alignment loop from the sensing matrix, to prevent some unwanted signals to survive this process. Please remember to remove any signal added to the sensing matrix after the need expires.
  • the hand-off 1f-3f is now completely automatic, and the reverse step, going back from 3f to 1f to recover from some unforeseen behaviour, is now implemented.

Towards the end of the shift, we tried once again to lock the green on the arms, but the green laser was still struggling to lock on the North and we were ultimately forced to give up. Noticeably, SNEB was still oscillating (Fig.2) during this attempt, and the SeaActivity alarm was still red. It is hard to say whether we could have been able to control the arms with the green without this issues.

At the end of the shift, we turned off the BPc and PSTAB loops from the North Arm ALS and we left the arms locked on the IR, with all the ALS and DRMI nodes in down.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×